OpenStudy (anonymous):

Given that f(x) = x2 − 3x + 3 and g(x) = the quantity of x minus one, over four, solve for f(g(x)) when x = 5.

OpenStudy (imstuck):

The trick to this is to work from the inside out. You have f(g(x)), so you will begin with the g(x)function, the one that's the "most" inside. Put 5 in for the x in the g function, like this:

OpenStudy (imstuck):

\[g(x)=\frac{ x-1 }{ 4 }-->\frac{ 5-1 }{4 }=\frac{ 4 }{4 }=1\]

OpenStudy (imstuck):

Now take the result from that, the 1, and replace the x in the f function with it, like this:

OpenStudy (imstuck):

\[f(1)=(1)^{2}-3(1)+3-->1-3+3=1\]So the whole thing equals 1.

OpenStudy (imstuck):

That's how you do those...work from the inside out.

OpenStudy (anonymous):

why do they have to make it so complicated why cant they just say that ! ugh !

OpenStudy (anonymous):

thank you so much !

OpenStudy (anonymous):

@IMstuck

OpenStudy (imstuck):

You're welcome!

OpenStudy (imstuck):

ty for the medal!

OpenStudy (anonymous):

:)